find the principle when :
final amount =4500 at 20 %per annum in 5 years

Answer by stanbon(75887)

find the principle when :
final amount =4500 at 20 %per annum in 5 years
------
A(t) = P(1+rt)
4500 = P(1+0.2*5)
4500 = P(1+1)
P = 4500/2
Priciple = $2250
===================
Cheers,
Stan H.
===================